a 50kg sack of rice, resting on the ground, is dragged by a 100 N forceat 30 degrees with the horizontal. if the friction is 8N, what is the acceleration of the sack
1
Expert's answer
2021-09-20T09:59:55-0400

Given:

"F=100\:\rm N"

"\theta=30^{\circ}"

"F_f=8\:\rm N"

"m=50\:\rm kg"


The Newton's second law says

"ma_x=F_x"

"ma_x=F\cos\theta-F_f"

"a_x=\frac{F\cos\theta-F_f}{m}"

"a_x=\frac{100\cos30^{\circ}-8}{50}=1.57\:\rm m/s^2"